1. A toroidal sample of magnetic material of susceptibility x = 2 × 10-2 is wound with 100 turns of wire carrying a current of 2 A. The toroid is 0.10 m long.

(a) Find the solenoidal current density.(3)

(b) Determine the magnetic field intensity H produced by the current.(3)

(c) Calculate μ, the magnetic permeability of the material.(3)

(d) Calculate the induced magnetization M in the material.(3)

(e) Calculate the magnetic field B resulting from the current and the magnetization of the material.(3)


1
Expert's answer
2021-06-30T10:19:15-0400

(a) The solenoidal current density depends on the turn density

"n=\frac{N}{l} \\

n = \frac{100}{0.10}=1000"

(b) The magnetic field intensity H produced by the current

"H= \frac{NI}{l} \\

H= \frac{100 \times 2}{0.10}=2000\;A/m"

(c) The magnetic permeability of the material

"\mu=\mu_0(1+χ)"

χ=susceptibility of the materials

"\mu=4 \pi \times 10^{-7}(1+2 \times 10^{-2}) \;H/m \\

\mu = 1.28 \times 10^{-6} \;H/m"

(d) The induced magnetization

"M= χH \\

M= 2 \times 10^{-2} \times 2000 = 40 \;A/m"

(e) The magnetic field B resulting from the current and the magnetization of the material

"B=\mu_0 nI \\

B= 4 \pi \times 10^{-7} \times 1000 \times 2 \\

B=2.51 \times 10^{-3} \; T"